Finding Slope and y-Intercept From an Equation.
For example:
-4x + 6y = 18

Answer by Alan3354(69443)

Finding Slope and y-Intercept From an Equation.
For example:
-4x + 6y = 18
------------
Solve for y
-4x + 6y = 18
y = (2/3)x + 3
The slope, m, is the coefficient of x
m = 2/3
-----
The y-int is the constant
b = 3
That's all, just solve for y.